Peminat mobil Toyota Innova diketahui mempunya marginal rate of substitution (MRS) dari space (luas interior) dan percepatan (acceleration) yang rendah. Sedang peminat mobil Ferarri mempunyai MRS yang tinggi. Gambarkan indifference curve dari kedua kelompok tersebut dan jelaskan.

Toyota Innova car enthusiasts are known to have a low marginal rate of substitution (MRS) of space (interior area) and acceleration (acceleration). Ferarri car enthusiasts have a high MRS. Draw the indifference curve of the two groups and explain.
Answer #1

Marginal rate of substitution refers to the amount of commodity an individual is willing to give up to consume one more unit of other commodity. In this case, MRS refers to the amount of space individual is willing to give up to get an additional unit of acceleration . MRS = dspace/dacceleration. In the question, it is given that Toyota enthusiasts have lower MRS, which means that they are willing to give up lower value of space for one more unit of acceleration as compared to the Ferrari enthusiasts. This implies that Toyota is a family car and space plays important role in utility while for Ferrari enthusiasts buy it for speed thrill and therefore, acceleration is more important to them than space.

It can be seen from the following diagram that for same change in acceleration, change in space is higher for ferrari enthusiasts as compared to Toyota enthusiasts, i.e. (aT1-aT0) = (aF1-aF0) = 1, then, (ST0-ST1) < (SF0-SF1).

Space Space I SE penari -IC Toyola acceleration acceleration